# Decentralization of Ethereum’s Builder Market

### v3

@@ 8 Conclusions
In this paper, we interrogate the causes and implications of builder centralization in Ethereum and challenge the community’s belief that it is acceptable. Our empirical study shows that the current blocker market has led to significant profit losses for validators. These significant losses could lead to PBS’s instability and MEV oracles’ inaccuracy and, if left uncontrolled, can lead to validator centralization. We attribute the loss to two factors: insufficient competition and inequality of block-building capacities, with the latter being the dominant reason. We further investigated the trust and incentive issues of the current MEV supply chain that contribute to builder centralization. To address these two problems, we propose a structural change to the current MEV supply chain and a solution based on this new structure.
@@ Appendix B Data Collection
Builder public keys.   We collect builder public keys from two data sources. First, builders often disclose their public keys in official documents. For example, Titan builder has disclosed 15 public keys in its official document [10]. Second, some builders leave their unique marks in the blocks’ “extra data” field. For example, Flashbots builder uses the mark ‘‘Illuminate Democratize Distribute’’. Given that such marks are unauthenticated and susceptible to impersonation, we ignore them unless they are used more than 100 times in the bids by the same builder public key.
Transaction sources.   We identified transactions originated from Telegram trading bots, including Maestro [55], Banana Gun [5], and Unibot [82], by searching for transactions that interacted with their official smart contracts.
Two channels, MEV Blocker and MEV-Share publish data [70, 33] about transactions submitted to them. Specifically, transactions submitted to MEV Blocker are publicly available on Dune [70], from which we extracted 22,819,437 transaction hashes that fall within our study’s range. We obtained 20,334,513 transactions submitted to MEV-Share from the public Flashbots dataset [33]. These transaction labels are sourced from their official datasets and we view them as ground truth.
Appendix C List of Searchers and Builders
In Tab. 3, we provide the list of searchers and their addresses mentioned in our study. In Tab. 4, we have counted the top 25 builders in the market, ranked by their market share from April 2023 to December 2023.
Table 3: The list of searchers
@@ Appendix C List of Searchers and Builders
Searcher: 0xA69b… 0xA69babEF1cA67A37Ffaf7a485DfFF3382056e78C
Searcher: 0x0087… 0x0087BB802D9C0e343F00510000729031ce00bf27
Searcher: $0$x280027dd00eE0050d3F9d168EFD6B40090009246
Searcher: 0x51c7… 0x51C72848c68a965f66FA7a88855F9f7784502a7F
Table 4: The percentages of blocks and MEV-Boost blocks built by each builder from April to December 2023.
Builder % of total blocks % of MEV-Boost blocks
Beaver 22.53 24.69
Rsync 20.75 22.74
builder0x69 12.65 13.86
Titan 12.35 13.53
Flashbots 12.24 13.42
BloXroute 2.14 2.34
f1b.io 1.66 1.81
Blocknative 1.40 1.53
eth-builder 0.77 0.84
BuildAI 0.73 0.80
Gambit Labs 0.64 0.70
Jetbldr 0.47 0.51
Manta 0.46 0.51
payload.de 0.39 0.42
boba-builder.com 0.27 0.30
Penguinbuild 0.23 0.25
Eden 0.22 0.24
lightspeedbuilder 0.21 0.23
0x3be…436 0.16 0.18
lokibuilder.xyz 0.12 0.13
bobTheBuilder.xyz 0.11 0.13
blockbeelder.com 0.11 0.12
finest artisanal blocks 0.09 0.10
It’s Free Real Estate 0.08 0.09
nfactorial.xyz 0.06 0.07
Appendix D Implications of Builder Centralization
@@ D.1 Efficiency of MEV-Boost Auctions
Analysis of inefficiency.   Fig. 4 shows that 24.9% of MEV-Boost auctions were not efficient. In an inefficient auction, the builder with the highest true value does not win the auction. There are two possible reasons. First, a builder can win the auction without possessing the highest true value by overbidding, known as block subsidization [36]. Second, builders may “shade” their bids (bid shading means strategically placing bids slightly below one’s true value); if the builder with the highest true value “shades too much,” it will lose.
@@
It is important to note that inefficiency does not necessarily disadvantage the validators. In fact, validators may receive more revenues in inefficient scenarios involving subsidization.
@@ D.2 Measuring Block-building Capacities
Figure 12: Violin graphs for the distribution of the CV across three MEV tiers, for top, middle, and tail builders. The CV measures the inequality of block-build capacity across builders.
@@ D.3 Implications of Uncompetitive Auctions
Figure 13: The distribution of the CIs of MEV-Boost auctions
Appendix E Reference Builder Algorithm
Flashbots has open-sourced a Rust implementation of the MEV-Boost builder [34], which serves as a reasonably performant reference for builder algorithms. We will briefly introduce their approach. The builder processes each bundle (where a single transaction is considered a bundle containing one transaction) by computing its effective gas price and total profit. Subsequently, the builder sorts all received bundles in descending order based on either effective gas price or total profit. Employing a greedy algorithm, the builder then appends bundles from this sorted list. 1 presents the pseudocode of their algorithm.
Listing 1: Flashbots block-building algorithm
# Compute the effective gas price and profit for a given bundle
# A single transaction is recognized as a bundle of size 1
def receive(bundle):
effective_gas_price, profit = compute_price_and_profit(bundle)
bundle.egp = effective_gas_price
bundle.profit = profit
return bundle
# Building a block
def build(bundles, sorting_key):
# Sort bundles by price or profit
if sorting_key == ’egp’:
bundles.sort(key=lambda b: b.egp)
else:
bundles.sort(key=lambda b: b.profit)
# Initialize
B = []
# Get the current state
state = get_current_state()
# Greedy algorithm
while not bundles.is_empty():
# Get the first bundle from the queue
bundle = bundles.pop(0)
# Simulate the execution of the bundle
success = simulate_bundle(state, bundle)
# Update with the successful bundle
if success:
B.append(bundle)
state.update(bundle)
# Compute payment based on the final state
payment = compute_payment(state)
# Append the payment to the block
B.append(payment)
# Return the valid block
return B
Appendix F What Caused Inequality in Block-building Capacities
F.1 Importance of Private Order Flows
To quantify the importance of private order flows for building valuable blocks, we compute the ratio of block values derived from private transactions. We first identify all private order flows using methods described in Sec. 3.1. Then, for each block, we compute the ratio between values extracted from private transactions (i.e., the total transaction priority fees [21] and direct transfers [32] to the block builder) to the total values extracted from all transactions in the block.
F.2 Pivotal Private Order Flow Providers
@@
Second, we can observe a positive trend from Fig. 8 that MEV Share and MEV Blocker have been increasingly pivotal since May 2023. These two providers positively impact decentralization because they have relatively clear requirements for accessing their order flows (the requirements of different providers will be discussed in Sec. 5.2).
@@
Although platforms like MEV-Share and MEV Blocker positively affect decentralization, the incentives of their upstream providers may not be aligned. For example, Mastro may prefer to send flows to builders directly for a better price than what MEV-Share and MEV Blocker can offer. It is surprising to see that upstream providers can significantly swing the significance of platforms.
Finding V: We identified five pivotal providers (MEV-Share, MEV Blocker, jaredfromsubway.eth, Banana Gun, Maestro) who had a sustained influence on the outcome of more than half of MEV-Boost auctions. A positive trend towards decentralization is observed as the pivotal level of the searcher jaredfromsubway.eth decreased, while the pivotal level of MEV-Share and MEV Blocker increased. However, their reliance on upstream providers like Maestro is alarming.
Case study for maestro.    To delve deeper into the relationship between these two trends, we plot Maestro’s contribution to the total daily value of transactions sent through MEV-Share and MEV Blocker during our study in Fig. 14. Surprisingly, during the time periods when MEV-Share and MEV Blocker were highly pivotal in Fig. 8, i.e., from May to November, Maestro’s contribution accounts for approximately 80%! This means Maestro used to be an important upstream of MEV-Share and MEV Blocker. In November 2023, Maestro stopped sending transactions to MEV-Share and MEV Blocker, causing the pivotal levels of MEV-Share and MEV Blocker to drop significantly.
Figure 14: Daily percentage of the value of the transactions originating from Maestro for both MEV-Share and MEV Blocker.
Figure 15: Weekly builder-specific pivotal levels for the top-3 builders for slots with winning bid values of no less than 0.5 ETH.
F.3 New Integration Pattern
To illustrate the new integration pattern between providers and builders, we consider the MEV significant in a slot if its winning bid value is no less than 0.5 ETH. Additionally, we focus only on the provider that is the sole pivotal provider in a slot, meaning this provider provides significant MEV to the builder, enabling the builder to win the auction. We apply the above adjustments to builder-specific pivotal levels, and the results are shown in Fig. 15. Compared to Fig. 9, we observe that Banana Gun and Titan, as well as jaredfromsubway.eth and Beaver, follow this integration pattern: a provider selectively shares valuable order flows with specific builders when MEV is significant.
F.4 Requirements of Providers
Tab. 5 summarizes the providers we identify and their requirements for accessing their private order flows.
@@ F.4 Requirements of Providers
Table 5: Collections of the private order flow providers and their requirements.
F.5 Minimal Cost for 1% Market Share
@@
The result is plotted in Fig. 16. Initially, this cost was nearly zero before May 2023, but it has steadily increased, surpassing 7.4 ETH by July 2024.It is important to recognize that the costs derived from this method only represent the lower bound. Predicting the minimum cost of winning a MEV-Boost auction is challenging, so builders need to pay more to ensure victory. Furthermore, builders may subsidize more blocks than required, as it is also difficult to precisely achieve a 1% market share. In reality, the subsidies new builders need to provide are substantially higher. For example, tbuilder [78], a new builder in the builder market, subsidized 6.8 ETH in total for 645 blocks in the week before March 21, 2024 [36].
@@
Appendix G Pseudocode of TEE Builder
The pseudocode for the builder running in TEEs with a refund mechanism is shown in Fig. 17.
[…]
$orders\leftarrow[]$ // Pending transactions.
$sources\leftarrow\{\}$ // Record the order flow providers.
// Receive order flows from providers.
[…]
$orders\leftarrow orders+[order]$
$sources[order]\leftarrow provider$
// Build best block and refund providers.
On building:
$v,B\leftarrow\textsc{build}(orders)$
$payments\leftarrow[]$
// For each provider, refund its marginal contribution,
 For every $order$ in $B$:
  $provider\leftarrow sources[order]$
  $v^{\prime},\_\leftarrow\textsc{build}(orders-[order])$
  $payment\leftarrow\textsc{transfer}(provider,v-v^{\prime})$
  $payments\leftarrow payments+[payment]$
// Remove all transactions included on chain.
 For every $order$ in $B$:
  $orders\leftarrow orders-[order]$
  Delete $sources[order]$
$B\leftarrow B+payments$
[…]
Figure 17: The pseudocode of TEE builder.

### v4

@@ 8 Conclusions
In this paper, we challenge the community’s belief that builder centralization is acceptable by empirically quantifying the significant profit losses incurred by validators in the current builder market. These significant losses could lead to PBS’s instability and MEV oracles’ inaccuracy and, if left uncontrolled, may lead to validator centralization. We attribute the loss to two factors: insufficient competition and inequality of block-building capacities, with the latter being the dominant reason. We further investigated the incentive issues of the current MEV supply centralization that contribute to the builder centralization and significant proposer loss. Finally, we identify the properties required to address these issues and hope these insights will benefit future work.
@@ Appendix B Data Collection
Builder public keys.   We collect builder public keys from three data sources. First, builders often disclose their public keys in official documents. For example, Titan builder has disclosed 15 public keys in its official document [76]. Second, some builders leave their unique marks in the blocks’ “extra data” field. For example, Flashbots builder uses the mark ‘‘Illuminate Democratize Distribute’’. We manually group builder public keys based on these marks and exclude the public keys which are known as impersonation [77].
Transaction sources.   We identified transactions originated from top Telegram trading bots, including Maestro [78], Banana Gun [79], and Unibot [80], by searching for transactions that interacted with their official smart contracts.
Two channels, MEV Blocker and MEV-Share publish data [81, 82] about transactions submitted to them. Specifically, transactions submitted to MEV Blocker are publicly available on Dune [81], from which we extracted 28,233,223 transaction hashes. We obtained 24,449,086 transactions submitted to MEV-Share from the public Flashbots dataset [82]. These transaction labels are sourced from their official datasets and we view them as ground truth.
Appendix C List of Searchers
In Tab. IV, we provide the list of searchers and their addresses mentioned in our study.
TABLE IV: The list of searchers
@@ Appendix C List of Searchers
SCP 0xA69babEF1cA67A37Ffaf7a485DfFF3382056e78C
0x56178a0d5F301bAf6CF3e1Cd53d9863437345Bf9
0xa57Bd00134B2850B2a1c55860c9e9ea100fDd6CF
0x4Cb18386e5d1F34dC6EEA834bf3534A970a3f8e7
0x4Cb18386e5d1F34dC6EEA834bf3534A970a3f8e7
0x5050e08626c499411B5D0E0b5AF0E83d3fD82EDF
0xFA103c21ea2DF71DFb92B0652F8B1D795e51cdEf
0x0DA9d9eCeA7235c999764e34F08499cA424c0177
WINTERMUTE 0x0087BB802D9C0e343F00510000729031ce00bf27
0x280027dd00eE0050d3F9d168EFD6B40090009246
0x51C72848c68a965f66FA7a88855F9f7784502a7F
Appendix D Efficiency and Block-building Capacities
@@ D.1 Efficiency of MEV-Boost Auctions
Analysis of inefficiency.   Fig. 4 shows that 24.9% of MEV-Boost auctions were not efficient. An auction is inefficient when the builder with the highest true value does not win. There are two possible reasons. First, a builder can win the auction without possessing the highest true value by overbidding, known as subsidization [70]. Second, builders may “shade” their bids (bid shading means strategically placing bids slightly below one’s true value); if the builder with the highest true value “shades too much,” it will lose.
@@
It is important to note that inefficiency does not necessarily harm validators. In fact, validators may receive more revenues in inefficient scenarios involving subsidization.
@@ D.2 Measuring Block-building Capacities
Figure 12: Violin graphs for the distribution of the CV across three MEV tiers, for top, middle, and tail builders. The CV measures the inequality of builders’ block-build capacity.
@@ D.3 Implications of Uncompetitive Auctions
Figure 13: Distribution of the CIs of MEV-Boost auctions
Appendix E Private Order Flows
E.1 Pivotal Private Order Flow Providers
@@
Second, we can observe a positive trend from Fig. 9 that MEV Share and MEV Blocker have been increasingly pivotal since May 2023. These two providers positively impact decentralization because they have relatively clear requirements for accessing their order flows (the requirements of different providers will be discussed next.)
@@
Although platforms like MEV-Share and MEV Blocker positively affect decentralization, the incentives of their upstream providers may not be aligned. For example, we observe that Maestro formed an integration with Beaver after ceasing to use these two platforms.
E.2 Requirements of Providers
Due to the lack of trustworthy fair exchange mechanisms, providers have to rely on an informal reputation system for self-protection, sending order flows only to builders who meet their requirements. To understand the access barrier caused by the trust crisis, we investigate different providers and their requirements for accessing order flows. Tab. V summarizes the providers we identify and their requirements for accessing their private order flows.
@@ E.2 Requirements of Providers
TABLE V: Collections of the private order flow providers and their requirements.
Subsidization.   New builders face a “chicken and egg” problem as they need access to private order flows to win auctions and gain market share, but the private order flow providers only serve builders with a decent market share. In practice, builders pay out of pocket to overbid in MEV-Boost auctions, a practice known as block subsidization [70], representing an entry barrier for new builders. We quantify the minimal cost for a new builder to meet the MEV Blocker’s requirements (Appx. E.3), which reflects the costs of building a reputation. The minimal cost continues to increase and can be up to 18 ETH. This cost represents a lower bound since we do not have the corresponding numbers for other providers. Note that although subsidization can help meet the requirements set by channels, it cannot ensure that searchers’ trust concerns are addressed, allowing builders to receive their order flows.
E.3 Minimal Cost for 1% Market Share
@@
The result is plotted in Fig. 14. Initially, this cost was nearly zero before May 2023, but it has steadily increased, surpassing 18 ETH by October 2024. It is important to recognize that the costs derived from this method only represent the lower bound. Predicting the minimum cost of winning a MEV-Boost auction is challenging, so builders need to pay more to ensure victory. Furthermore, builders may subsidize more blocks than required, as it is also difficult to precisely achieve a 1% market share.
@@
Appendix F Limitations of Our Approach and Data
True value.   We compute the true values of bids using a similar approach as that in MEV analysis tools [41, 42] and previous works [43, 44]. This approach has a few limitations. First, it relies on the knowledge of the set of addresses controlled by each builder. Most builders use well-known addresses to receive payments from searchers and make payments to proposers, but it is possible that some builders use less publicized addresses. We cross-validated our builder address dataset with other sources to ensure completeness (see Sec. 3.1), but if we missed a builder’s address, the computation of extracted value (and hence true value) would be inaccurate.
Second, this approach may neglect off-chain profits, such as off-chain payments received from searchers. Off-chain payments are more likely to happen between integrated builder-searcher pairs. We might underestimate the true value of a block due to unknown off-chain payments, which could lead to the misclassification of an efficient auction as inefficient due to subsidization. However, in our study of evaluating competitiveness and efficiency, we did not observe any excessively large subsidies. Furthermore, underestimating the winner’s true value does not affect our assessment of competitiveness. We might misclassify an uncompetitive auction as competitive if we underestimate the second-highest true value. However, such underestimation should be uncommon, as we observe that most builders bid around our computed true value. It would be irrational for builders not to bid higher to win the auction if they actually have a higher true value.
Third, this approach may overestimate builders’ profits because builders might refund providers for their private order flows. Such refunds may occur through private, off-chain deals and are therefore not captured. Overestimating the winner’s true value does not impact the competitiveness calculation; however, overestimating the second-highest true value could lead us to misclassify a competitive auction as uncompetitive. However, our study indicates that this kind of integration between top builders and providers began on a large scale starting in 2024, after the time range of our evaluation of competitiveness and efficiency.
Missing pivotal providers.   Our results in Sec. 5.1 might omit some pivotal providers since the providers we can identify are a subset of all providers. For example, BloXroute (a widely used channel) does not publish the list of transactions they processed. Thus, we can’t evaluate its pivotal level. This limitation is common in existing studies [18, 19]. However, there will not be false positives (classification of a non-pivotal provider as pivotal) because we identify a given transaction’s provider based either on the official dataset from MEV-Share and MEV Blocker or on the specific searchers’ addresses (see Sec. 3). Thus, the significance of providers will only be underestimated.